OpenStudy (anonymous):

a boat can travel 8 miles downstream in the same amoun tof time it can travel 6 miles upstream. the speed of the boat in still water is 14 mph. find the speed of the current?

hero (hero):

d = rt r = b±w d = (b±w)t 8 = (14+w)t 8 = (14-w)t 8 = 14t + wt 8 = 14t - wt wt = 8 - 14t wt = 14t - 8 wt = wt 8-14t = 14t - 8 16 = 28t 16/28 = t 4/7 = t 4/7 hrs 4/7 of 60 min ≈ 34 min t ≈ 34 8 = (14-w)4/7 8 = 8 - 34w 34w = 8 - 8 34w = 0 w = 0/34 w = 0 Thus the speed of the current is zero, which makes sense because that's the only possible way a boat can travel at the same speed up and downstream in the same time.

hero (hero):

@shivraj